OpenStudy (spareb665):

Solve the equation 2^3x+1=5^-x and round your answer to two digits after the decimal sign

OpenStudy (anonymous):

2^(3x+1) =5^(-x) (3x+1)Ln(2) = -xLn(5) (3x+1)/(-x)=(Ln(5)/Ln(2) -3-(1/x)=(Ln(5)/Ln(2) -(1/x)=(Ln(5)/Ln(2)+3 x=-1/((Ln(5)/Ln(2)-3)